OpenStudy (anonymous):

Write an equation that describes the following variation. V varies directly with the cube of r. V=28pi when r= 4.

OpenStudy (anonymous):

V=k*r^3 ==> k=28 pi /4= 7 pi ==> \[ V=7\pi r^{3}\]

OpenStudy (anonymous):